The Perimeter of a rectangle is 6x+6 yds. If the width is x yards, find a binomial that represents the length.
Answer by rapaljer(4671)

Let x = width
Let L = Length
P = 6x+6

2x+2L= P
2x+2L= 6x+6, you must solve for L.

Subtract 2x from each side.
2x-2x+2L = 6x+6-2x
2L= 4x+6

Divide both sides by 2.
L= 2x+3

Dr. Robert J. Rapalje

P.S. Please visit my website at "www2.seminolestate.edu/rrapalje". It's a pretty good website with lots of good math materials and a nice JOKE page!! I need the hits!!